# UTMath: Math Evaluation with Unit Test via Reasoning-to-Coding Thoughts

### v1

@@ 3.3 Evaluation Metrics
We adopt the metric pass@$k$ to evaluate the performance of LLMs. The metric pass@$k$ is a classic metric in code generation, where k code samples are generated per problem, a problem is considered solved if any sample passes the unit tests, and the total fraction of problems solved is reported. We use the stable method of calculation proposed by (Chen et al., 2021):
@@
3.4 Dataset Statistics
To gain a deeper understanding of the composition of the UTMath Benchmark, we analyzed the potential fields each problem might pertain to and identified nine mathematical fields: Number Theory, Graph Theory, Group Theory, Game Theory, Discrete Mathematics, Combinatorial Mathematics, Geometry and Topology, Polynomial and Series Expansions, Special Numbers, Formal Languages. Using GPT-4o, we allocated each problem to these fields and obtained the distribution shown in Table 1.
Table 1: Categories and distribution of problems
Category Number
@@ 3.4 Dataset Statistics
Total Problems 1053
Table 2: Pass Rate and Average Run Time of LLMs on UTMath
@@ 4 Reasoning-to-Coding of Thoughts
Compared to evaluation methods that solely rely on checking if the results generated by LLMs are identical, without assessing the reasoning process, we propose a method based on evaluating mathematical reasoning processes through code implementation.
Initially, we adopted the Program of Thought (PoT) method, where the LLM had to perform reasoning and implement it in one step. However, we noticed that the LLMs often resorted to simpler algorithms, which led to high time complexity or even failure with more complex problems due to limited reasoning depth. To improve this, we explored the Reasoning-to-Coding of Thoughts framework, which separates reasoning and implementation into different steps.
In the first round, LLM focuses only on reasoning about the problem. Compared to PoT, where reasoning is described broadly, dedicating an entire round to reasoning allows the LLM to generate a step-by-step, detailed logical reasoning chain, including mathematical theorems, formulas, and properties used. This deeper reasoning approach reduces the likelihood of errors and facilitates the creation of less complex algorithms.
In the second round, the LLM implements the reasoning process generated in the first round.This gives us new perspectives beyond the final answer when evaluating reasoning abilities. By converting the computation process into code, we avoid errors caused by low computational skills, allowing us to gain more accurate and genuine insights into the LLM’s reasoning abilities. Additionally, assessing the pass rate and runtime of the code provides a method to evaluate reasoning strategies, where better approaches should have higher accuracy and lower time complexity.
By dividing the process into two steps, The detailed reasoning in the first round ensures that the code implementation in the second round closely aligns with the reasoning process, reducing mismatches between reasoning and implementation.
@@ 5.1 Experimental Setup
Here, we consider the closed-source models, i.e., GPT-3.5-Turbo/GPT-4o from OpenAI OpenAI (2024), Claude-3.5-Sonnet Claude (2024), Gemini-1.5-Pro Reid et al. (2024), as well as the open-source models, i.e., LLaMA-3.1 Dubey et al. (2024), Qwen2.5 Qwen (2024a), Qwen2.5-Math Qwen (2024b), DeepSeek-V2.5 Bi et al. (2024). The metric pass@$1$ is calculated as the average result over 5 run times. We run all evaluations in a laptop with CPU Intel(R) Core(TM) i7-10750H CPU @ 2.60GHz.
@@ 5.2 Evaluation on UTMath
Here we evaluate both open-source and closed-source models using RCoT and PoT in Tab. 2. The experimental results shows that all tested models performed poorly on our benchmark. The best model, GPT-4o, only solves 26.93% problem in our benchmark. Since our problems are sourced from the OEIS, they consist of sequences and solutions proposed by various mathematicians in the context of cutting-edge research. This suggests that our benchmark is challenging enough to help guide future directions for improving LLMs.
Compared to PoT, our method RCoT demonstrates superiority in two aspects. First, prompting with RCoT achieves higher pass@$5$ performance across 6 LLMs, with the best results observed on GPT-4o. Second, the solutions generated by RCoT demonstrate more efficient performance, particularly Qwen2.5-72B, where the RCoT approach achieves an efficiency improvement of over 36.42% compared to PoT, as shown in Tab. 2 and Fig. 2. It indicates that, with RCoT prompting, the model engages in deeper reasoning, significantly reducing solution complexity and enhancing solution performance. However, some models experienced a decrease in pass@$1$ with RCoT . Gemini-1.5-Pro, GPT-3.5-Turbo, and Qwen2.5-72B’s accuracies slightly dropped. Qwen2.5-72B’s dropped on pass@$1$ but increased on pass@$5$. It indicates that RCoT brings more room in multiple inference times. The poorer performance may be due to the fact that attempting to devise a more efficient solution often involves a higher level of difficulty, and these models are more prone to errors when reasoning through more complex solutions.
5.3 The Effectiveness of Hard Test Cases
As we mentioned in Sec. 3.2, each sequence in the OEIS lists only the initial terms, which we refer to as “easy test cases”. To investigate the model’s ability to handle challenging cases, we evaluated whether it could predict values that appear later (i.e., $10^{6}$) in a sequence. These later values are typically underrepresented in pre-training data and often require more computation time and a more precise implementation to retrieve accurately. The experimental results, which depicted in Tab. 3, reveal that the model’s performance drops significantly when handling these hard cases. This indicates that introducing these cases can prevent simple solutions from passing all the easy cases, thereby filtering for more advanced solutions. Our benchmark provides a better measure of the reasoning abilities across different models.
Table 3: Performance of different models on easy and hard test cases. Divide into open-source and closed-source models, and sort each model in descending order based on the performance on Easy & Hard cases.
@@ 5.3 The Effectiveness of Hard Test Cases
Table 4: Performance on Different Problem Categories.(%) Categories are represented by abbreviations. NT: Number Theory; T.: Theory; DM: Discrete Mathematics; CM: Combinatorial Mathematics; GT: Geometry and Topology; PSE: Polynomial and Series Expansions; SN: Special Numbers; FL: Formal Languages;
@@
GPT-4o 43.9 2.78 11.69 38.23 24.94 3.43 16.42 33.89 42.5 26.93
Claude-3.5-Sonnet 33.58 1.52 8.0 29.49 12.91 5.43 11.52 26.62 20.36 19.11
Gemini-1.5-Pro 22.64 1.27 5.23 19.75 8.35 2.29 8.61 14.01 15.36 12.35
GPT-3.5-Turbo 13.08 0.0 1.85 11.39 3.29 0.29 2.78 10.96 8.93 6.82
@@
Qwen2.5-72B 36.86 2.53 12.0 30.63 15.95 6.0 18.15 29.43 24.29 22.17
DeepSeek-V2.5 38.24 1.27 8.92 33.16 17.34 2.29 12.45 31.08 20.0 21.63
Qwen2.5-Math-72B 35.35 1.27 8.62 28.73 14.81 4.0 17.48 28.15 20.0 20.53
LLaMA-3.1-405B 29.56 0.76 4.92 25.44 9.62 2.0 9.54 22.55 21.43 16.09
@@ 5.4 Scaling of the Inference Times
We compared the performance difference between running the LLMs five times and reported the metric of pass@$k$. As shown in Fig. 4, all models improved their performance with an increasing number of inference times. For Qwen2.5-72B, RCoT was slightly weaker than PoT in the first inference but quickly approached and surpassed PoT in subsequent run times. For GPT-3.5-Turbo and Gemini-1.5-Pro, PoT consistently outperformed RCoT. We observed that with an increasing number of inference time, RCoT consistently demonstrated a growing advantage in performance across almost all models, except for GPT-3.5. This suggests that RCoT may perform better in models with stronger reasoning capabilities.
@@
5.5 Importance of the Reasoning Step
GPT-4o has the best performance, while we are unclear whether this was due to its superior reasoning or coding ability. To investigate further, we tested using GPT-4o for the reasoning step while other models perform the coding step based on the reasoning result from GPT-4o. As depicted in Fig.5, the results showed that the performance of models, except for Qwen2.5-72B, increased significantly when implementing coding based on GPT-4o’s reasoning output, suggesting that the reasoning quality is important and GPT-4o does produce higher-quality reasoning results.
@@ 5.6 Performance on Different Categories
Our benchmark comprehensively evaluates the LLMs’ ability across various categories of math problems. We analyzed the differences in performance in terms of categories. GPT-4o continued to demonstrate a clear advantage across most tasks, achieving the best performance in five domains. Notably, it outperformed the second-best model by 7.6% in combinatorial mathematics and by 18.21% in formal languages. In contrast, as an open-source model, Qwen2.5-72B secured the top spot in the other three domains. All models performed very poorly in the categories of Graph Theory, Group Theory, and Geometry and Topology, with accuracy rates below 12%, highlighting the need for further exploration in these areas.
@@ Appendix A An Example Sequence in OEIS
An example sequence is shown in Fig. 6
Figure 6: Sequence A001776 in OEIS.
Appendix B Rulers for Data Cleaning
@@
    Problem Description: The sequence is too difficult, requiring extensive background knowledge, or only a limited number of terms are found.
    Filtering Method: Remove sequences with keywords containing ‘hard’, ‘fin’ (finite).
@@
    Problem Description: The sequence is hard to generate with a program.
    Filtering Method: Check if it contains program, formula, or Mathematica fields in the sequence’s json data.
@@
    Problem Description: The sequence is too simple with an explicit recurrence or closed formula.
    Filtering Method: Search if the description includes ‘a(n) =’.
@@
    Problem Description: Solving the sequence requires information from other OEIS sequences.
    Filtering Method: Search if the sequence’s description contains the AID of other sequences (‘A’ + six-digit number with leading zeros).
@@
    Problem Description: The sequence is decimal expansion of a certain number.
    Filtering Method: Search if the description includes ‘decimal’.
@@
    Problem Description: The sequence consists of repetitions or a constant value.
    Filtering Method: Search if the description includes both ’repeat’ and ’period’ or ’constant sequence’.
@@
    Problem Description: The description is too vague.
    Filtering Method: Search if the description includes ‘related to’.
@@
    Problem Description: Another version of a concept.
    Filtering Method: Search if the title includes ‘another version’, ‘second kind’, etc.
@@
    Problem Description: The sequence is formed by taking mod of a constant.
    Filtering Method: Search if the description includes ‘module’.
@@
    Problem Description: The values in the sequence are too large, which might cause LLM tokenization errors.
    Filtering Method: Check if any term’s length exceeds 18 digits(i.e., greater than $181e$ 18), remove it.
@@
    Problem Description: Coefficient triangles or ‘read by row’ topics.
    Filtering Method: Search if the sequence’s description includes ‘read by row’, ‘triangle of coefficient’.
@@
    Problem Description: The description is too short, either purely implementation or lacks necessary information.
    Filtering Method: Check if the title length is below 5.
@@
    Problem Description: More like a reasoning puzzle.
    Filtering Method: Use GPT-4o to judge, with the prompt outlined in the prompts section of the appendix 9.
@@
    Problem Description: Non-mathematical topics.
    Filtering Method: Use GPT-4o to judge, with the prompt outlined in the prompts section of the appendix 11.

### v2

@@ 3.3 Evaluation Metrics
We adopt the metric pass@$k$ to evaluate the performance of LLMs. The metric pass@$k$ is a classic metric in code generation, where a problem is solved if any of the k generated samples passes the unit tests. We use the stable method of calculation proposed by (Chen et al., 2021):
@@
Category # of Problems
@@
Total 1053
Table 2: Categories and distribution of problems.
@@
Table 3: Pass Rate and Average Run Time of LLMs on UTMath. We listed the performance of eight large models by the PoT or the RCoT methods across a range of metrics. For o1-mini and o1-preview only Pass@1 data is currently available due to resource constraints. The average run time is calculated based on the problems solved by both the PoT and the RCoT methods. The efficiency is calculated as: (Avg.Runtime(PoT) - Avg.Runtime(RCoT)) / Avg.Runtime(RCoT). Two qualitative cases are shown in Appendix D.
3.4 Dataset Statistics
The main statistics of UTMath are shown in Tab. 1. To gain a deeper understanding of the composition of the UTMath Benchmark, we identified nine mathematical fields and used GPT-4o to categorize each problem to these fields as shown in Tab. 2. Our analysis reveals that only 10 out of 1,053 problems have no references. The reference years span from 1950 to 2024, with the maximum number of references exceeding 6,000. These findings underscore the cutting-edge nature of our benchmark. More details can be found in Appendix B.
@@ 4 Reasoning-to-Coding of Thoughts
Compared to methods that simply check whether the outputs generated by LLMs are identical, the code-based evaluation approach enables more accurate assessment by using multiple test cases. It provides additional evaluation metrics, such as runtime, and allows for the observation of the reliability of LLMs.
Initially, we adopted the Program of Thought (PoT) method, where the LLM had to perform reasoning and implement it in one step. However, we noticed that the LLMs often resorted to simpler algorithms, which led to high time complexity or even failure with more complex problems due to limited reasoning depth. To improve this, we explored the Reasoning-to-Coding of Thoughts (RCoT) framework, which separates reasoning and implementation into different steps.
In the first round, LLM focuses only on reasoning about the problem. Compared to PoT, RCoT dedicate an entire round to reasoning allowing the LLM to generate a step-by-step, detailed logical reasoning chain, including mathematical theorems, formulas, and properties used. This deeper reasoning approach facilitates the creation of more efficient algorithms with lower time complexity.
In the second round, RCoT requires the LLM to implement the reasoning process generated in the first round. By converting the computation process into code, we can introduce new metrics beyond the final pass rate. The solution generated by the LLM can be evaluated based on its runtime during testing, which indirectly reflects the solution’s time complexity. Additionally, this approach avoids errors caused by the limited computational capabilities of large models, allowing for more accurate and genuine insights into the LLM’s reasoning abilities. Clearly, the stronger the reasoning capabilities of the large model, the higher the overall pass rate and the lower the time complexity, which is reflected in the shorter runtime of the generated solution.
We present qualitative cases in Appendix D, where GPT-4o solves problems from UTMath using PoT and RCoT, respectively. The case studies show that, with RCoT prompting, the model engages in deeper reasoning, significantly reducing solution complexity.
@@ 5.1 Experimental Setup
Here, we consider the closed-source models, i.e., GPT-3.5-Turbo, GPT-4o, o1-mini and o1-preview from OpenAI OpenAI (2024), Claude-3.5-Sonnet Claude (2024), Gemini-1.5-Pro Reid et al. (2024), as well as the open-source models, i.e., LLaMA-3.1 Dubey et al. (2024), Qwen2.5 Qwen (2024a), Qwen2.5-Math Qwen (2024b), DeepSeek-V2.5 Bi et al. (2024). The metric pass@$1$ is calculated as the average result over 5 run times. We run all evaluations in a laptop with CPU Intel(R) Core(TM) i7-10750H CPU @ 2.60GHz.
@@ 5.2 Evaluation on UTMath
Here we evaluate both open-source and closed-source models using RCoT and PoT in Tab. 3. The experimental results shows that all tested models performed poorly on our benchmark. The best model, o1-mini, only solves 32.57% problem in our benchmark, followed by o1-preview at 27.16% and GPT-4o at 26.93%. Since our problems are sourced from the OEIS, they consist of sequences and solutions proposed by various mathematicians in the context of cutting-edge research. This suggests that our benchmark is challenging enough to help guide future directions for improving LLMs.
Compared to PoT, our method RCoT demonstrates superiority in two aspects. First, prompting with RCoT achieves higher pass@$5$ performance across 7 LLMs, with the best results observed on GPT-4o. Second, the solutions generated by RCoT for all LLMs demonstrate more efficient performance, particularly Qwen2.5-72B, where the RCoT approach achieves an efficiency improvement of over 36.42% compared to PoT, as shown in Tab. 3 and Fig. 2. It indicates that, RCoT prompting enables the model to engage in deeper reasoning, significantly reducing solution complexity and enhancing solution performance. However, some models experienced a decrease in pass@$1$ with RCoT . Specifically, the accuracies of Gemini-1.5-Pro, GPT-3.5-Turbo, and Qwen2.5-72B slightly dropped. Notably, while Gemini-1.5-Pro and Qwen2.5-72B experienced a drop in pass@$1$, their pass@$5$ performance improved. It indicates that RCoT brings more room in multiple inference times. The observed decrease in performance may stem from the fact that formulating more efficient solutions often requires higher-level reasoning, which can increase the difficulty of the task and make these models more susceptible to errors when attempting more sophisticated solutions.
@@
Table 4: Performance (%) of different models on easy and hard test cases. Easy cases: The initial terms in OEIS. Hard cases: mined hard test cases (§ 3.2).
@@
o1-mini 52.83 7.59 15.38 42.41 32.27 7.14 23.84 40.13 37.50 32.57
o1-preview 47.17 6.33 13.85 34.17 25.32 2.86 23.18 29.94 33.93 27.16
GPT-4o 43.90 2.78 11.69 38.23 24.94 3.43 16.42 33.89 42.50 26.93
Gemini-1.5-Pro 31.70 1.27 8.92 27.47 15.19 5.71 15.23 27.39 17.86 19.43
Claude-3.5-Sonnet 33.58 1.52 8.00 29.49 12.91 5.43 11.52 26.62 20.36 19.11
GPT-3.5-Turbo 13.08 0.00 1.85 11.39 3.29 0.29 2.78 10.96 8.93 6.82
@@
Qwen2.5-72B 36.86 2.53 12.00 30.63 15.95 6.00 18.15 29.43 24.29 22.17
DeepSeek-V2.5 38.24 1.27 8.92 33.16 17.34 2.29 12.45 31.08 20.00 21.63
Qwen2.5-Math-72B 35.35 1.27 8.62 28.73 14.81 4.00 17.48 28.15 20.00 20.53
LLaMA-3.1-405B 29.56 0.76 4.92 25.44 9.62 2.00 9.54 22.55 21.43 16.09
Table 5: Performance (%) on different problem categories. Categories are represented by abbreviations. NT: Number Theory; T.: Theory; DM: Discrete Mathematics; CM: Combinatorial Mathematics; GT: Geometry and Topology; PSE: Polynomial and Series Expansions; SN: Special Numbers; FL: Formal Languages.
5.3 The Effectiveness of Hard Test Cases
As we mentioned in § 3.2, each sequence in the OEIS lists only the initial terms, which we refer to as “easy test cases”. To investigate the model’s ability to handle challenging cases, we evaluated whether it could predict values that appear later (i.e., $10^{6}$) in a sequence. These later values are typically underrepresented in pre-training data and often require more computation time and a more precise implementation to retrieve accurately. The experimental results, which depicted in Tab. 4, reveal that the model’s performance drops significantly when handling these hard cases. This indicates that introducing these cases can prevent simple solutions from passing all the test cases, thereby filtering for more advanced solutions.
@@ 5.4 Scaling of the Inference Times
We compared the performance difference between running the LLMs five times and reported the metric of pass@$k$. As shown in Fig. 4, all models improved their performance with an increasing number of inference times. For Qwen2.5-72B and Gemini-1.5-Pro, RCoT was slightly weaker than PoT in pass@$1$ but quickly approached and surpassed PoT in subsequent run times. We observed that with an increasing number of inference time, RCoT consistently demonstrated a growing advantage in performance across almost all models, except for GPT-3.5. However, it is worth noting that GPT-3.5 exhibited the lowest pass rate. This suggests that RCoT may perform better in models with stronger reasoning capabilities.
5.5 Importance of the Reasoning Step
GPT-4o has the best performance, while we are unclear whether this was due to its superior reasoning or coding ability. To investigate further, we tested using GPT-4o for the reasoning step while other models perform the coding step based on the reasoning result from GPT-4o. As depicted in Fig.5, the results showed that the performance of models increased significantly when implementing coding based on GPT-4o’s reasoning output, suggesting that the reasoning quality is important and GPT-4o does produce higher-quality reasoning results.
@@ 5.6 Performance on Different Categories
Our benchmark comprehensively evaluates the LLMs’ ability across various categories of math problems. GPT-4o achieved the highest score in the formal language domain, while o1-mini achieved the best scores in the remaining eight domains. All models performed very poorly in the categories of Graph Theory, Group Theory, and Geometry and Topology, with accuracy rates below 16%, highlighting the need for further exploration in these areas.
6 Conclusion
In this work, we investigate how to more accurately and effectively evaluate the mathematical reasoning capabilities of LLMs. We propose a cutting-edge benchmark, UTMath, which comprises 1,053 problems spanning nine mathematical domains, with an average of 68 test cases per problem. This benchmark presents significant challenges: o1-mini, the best-performing model, successfully solves only 32.57% of the problems, followed by o1- preview at 27.16%, and GPT-4o at 26.93%. Additionally, we introduce RCoT (Reasoning-to-Coding of Thought). Our study finds that, compared to PoT (Program-of-Thought), RCoT significantly improves the algorithmic efficiency and pass rates of most models by facilitating deeper reasoning steps prior to code generation. Overall, this research contributes to a deeper understanding of the current capabilities of LLMs in mathematical reasoning and lays the groundwork for the development of more advanced models in the future.
Limitation
The primary limitation of UTMath lies in the evaluation metrics: the performance of the evaluation machine affects the runtime of the generated code, making the absolute numerical results incomparable across different machines. We utilized an i7-10750H processor to execute the reference solutions and conduct evaluations, and we recommend using the same machine for testing and replication. There are two main limitations of RCoT. First, we only installed a set of common packages, such as sympy, in the standard testing environment. This avoids allowing LLMs to call highly integrated packages while also preventing the generation of potentially harmful code that could damage the evaluation system. Second, while our experiments demonstrate the critical role of reasoning quality in determining success rates, we have not further explored methods for enhancing reasoning quality, which remains an area for future investigation.
Ethics Statements
The UTMath Benchmark is designed to advance the evaluation of mathematical reasoning in LLMs. We recognize the potential ethical concerns associated with this work, particularly the risk of data misuse. To mitigate this, we strictly adhere to usage guidelines and licensing terms for the UTMath-Train dataset, which is intended solely for academic and research purposes. While the UTMath Benchmark evaluates model performance in terms of accuracy and generality, automated evaluations may introduce biases due to the nature of the datasets and evaluation algorithms. Additionally, while UTMath covers a wide range of mathematical domains, it may not fully represent diverse cultural or educational perspectives. We encourage further development of benchmarks that incorporate a broader array of reasoning styles to ensure more inclusive evaluations. By releasing UTMath, we aim to foster responsible AI development, promoting better, more generalizable mathematical reasoning systems.
@@ Appendix A An Example Sequence in OEIS
Figure 6: Sequence A007369 in OEIS. Its description: "Numbers n such that sigma(x) = n has no solution." (Clearly, without specific background knowledge, we cannot fully understand what the function sigma() represents, which is one of the reasons we perform standardization. §B.2) Next, OEIS shows the first 67 terms of this sequence, which we classify as easy cases. Below that, additional metadata is provided, including comments, references, links, formulas, examples, programs, author, status, and more. It is evident that this sequence has garnered significant attention from researchers, reflecting the Cutting-Edge difficulty of our benchmark. We used the Mathematica program included in the metadata to generate Hard cases, with detailed procedures provided in § 3.2. As a scientific database, each sequence submitted to OEIS undergoes a review process, and the status "approve" indicates that the sequence has been validated and approved by OEIS administrators.
Appendix B Dataset Construction Details
This section primarily presents some details on the construction of UTMath.  B.1 discusses the issues encountered when observing data crawled from OEIS, along with the corresponding cleaning rules. UTMath applies all 14 rules. Additionally, we crawled all sequences from OEIS and, for convenience, applied only the first 12 rules to create UTMath_Train, which contains over 70k sequences.  B.2 outlines the process followed for standardizing the descriptions of problems in UTMath, while  B.3 explains the referencing of sequences within UTMath, highlighting both the Cutting-Edge difficulty level of UTMath and its scalability.
B.1 Rules for Data Cleaning
@@
    Issues: The sequence is too difficult, requiring extensive background knowledge, or only a limited number of terms are found.
    Method: Remove sequences with keywords containing ‘hard’, ‘fin’ (finite).
@@
    Issues: The sequence is hard to generate with a program.
    Method: Check if it contains program, formula, or Mathematica fields in the sequence’s json data.
@@
    Issues: The sequence is too simple with an explicit recurrence or closed formula.
    Method: Search if the description includes ‘a(n) =’.
@@
    Issues: Solving the sequence requires information from other OEIS sequences.
    Method: Search if the sequence’s description contains the AID of other sequences (‘A’ + six-digit number with leading zeros).
@@
    Issues: The sequence is decimal expansion of a certain number.
    Method: Search if the description includes ‘decimal’.
@@
    Issues: The sequence consists of repetitions or a constant value.
    Method: Search if the description includes both ’repeat’ and ’period’ or ’constant sequence’.
@@
    Issues: The description is too vague.
    Method: Search if the description includes ‘related to’.
@@
    Issues: Another version of a concept.
    Method: Search if the title includes ‘another version’, ‘second kind’, etc.
@@
    Issues: The sequence is formed by taking mod of a constant.
    Method: Search if the description includes ‘module’.
@@
    Issues: The values in the sequence are too large, which might cause LLM tokenization errors.
    Method: Check if any term’s length exceeds 18 digits(i.e., greater than $181e$ 18), remove it.
@@
    Issues: Coefficient triangles or ‘read by row’ topics.
    Method: Search if the sequence’s description includes ‘read by row’, ‘triangle of coefficient’.
@@
    Issues: The description is too short, either purely implementation or lacks necessary information.
    Method: Check if the title length is below 5.
@@
    Issues: More like a reasoning puzzle.
    Method: Use GPT-4o to judge, with the prompt outlined in the Appendix C.
@@
    Issues: Non-mathematical topics.
    Method: Use GPT-4o to judge, with the prompt outlined in the Appendix C.
B.2 Standardization of Problems’ Description
Figure 7: Comparison between original and standardized problem description. The standardized version includes hints and explains the specific meaning of $\sigma(x)$.
B.3 Dataset Statistics
To demonstrate that our benchmark is of cutting-edge level, we have analyzed the distribution of the publication years and the number of references included in the problems of the benchmark as shown in Fig. 8. Additionally, OEIS is a dynamic database. Over the past five years, more than 35,000 sequences in UTMath_Train have been further researched, and over 2,000 new sequences have been added. This ongoing development makes it possible to continuously update UTMath_Train and UTMath, helping to address the challenges posed by data leakage.
Figure 8: Distribution of references in UTMath.